x+10/x = 18/12 solve for x
Answer by harpazo(655)

(x+10)/x = 18/12


Cross-multiply to solve for x.


18x = 12(x + 10)


18x = 12x + 120


18x - 12x = 120


6x = 120


x = 120/6


x = 20